Heyden Company has fixed costs of $381,000. The unit selling price, variable cost per unit, and contribution margin per unit for the company's two products are provided below.

Product Selling Price Variable Cost per Unit Contribution Margin per Unit
Model 94 $110 $60 $50
Model 81 170 120 50

The sales mix for products Model 94 and Model 81 is 70% and 30%, respectively. Determine the break-even point in units of Model 94 and Model 81 of the overall (total) product, E. If required, round your answers to the nearest whole number.

a. Product Model 94  units

b. Product Model 81  units
Answer #1

Weighted average contribution margin=Respective contribution margin*Respective sales mix

=(50*0.7)+(50*0.3)

=$50

Breakeven=Fixed expenses/Weighted average contribution margin

=381,000/50

=7620 units

a. Product Model 94  units=(7620*70%)=5334 units

b. Product Model 81  units=(7620*30%)=2286 units